how do you solve the equation down below?
x+y=8
x-y=2

Answer by checkley79(3341)

x+y=8
x-y=2 add.
--------
2x=10
x=10/2
x=5 ans.
5+y=8
y=8-5
y=3 ans.
Proof:
5-3=2
2=2